# Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation

### v1

@@ B.1 Proof of the identification of latent variables
Proof.
We begin with the matched marginal distribution $p_{\mathbf{x}|\mathbf{e},\mathbf{y}}$ to bridge the relation between $\mathbf{z}$ and $\hat{\mathbf{z}}$. Suppose that $\hat{g}:\mathbf{z}\rightarrow\mathcal{X}$ is an invertible estimated generating function, we have Equation (10).
$\forall\mathbf{e}\in\mathcal{(}E),\quad p_{\hat{\mathbf{x}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{x}|\mathbf{e},\mathbf{y}}\Longleftrightarrow p_{\hat{g}(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{g(\mathbf{z})|\mathbf{e},\mathbf{y}}.$ (10)
Sequentially, by using the change of variables formula, we can further obtain Equation (11)
$p_{\hat{g}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}=p_{g(\mathbf{z}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{g^{-1}\circ g(\hat{\mathbf{z}})|\mathbf{e}}|\mathbf{J}_{g^{-1}}|=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{g^{-1}}|\Longleftrightarrow p_{h(\hat{\mathbf{z}})|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}},$ (11)
where $h:=g^{-1}\circ g$ is the transformation between the ground-true and the estimated latent variables, respectively. $\mathbf{J}_{g^{-1}}$ denotes the absolute value of Jacobian matrix determinant of $g^{-1}$. Since we assume that $g$ and $\hat{g}$ are invertible, $|\mathbf{J}_{g^{-1}}|\neq 0$ and $h$ is also invertible.
According to A2 (conditional independent assumption), we can have Equation (12).
$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{z_{i}|\mathbf{e},\mathbf{y}}(z_{i}|\mathbf{e},\mathbf{y});\quad p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})=\prod_{i=1}^{n}p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}(\hat{z}_{i}|\mathbf{e},\mathbf{y}).$ (12)
For convenience, we take logarithm on both sides of Equation (12) and further let $q_{i}:=\log p_{z_{i}|\mathbf{e},\mathbf{y}},\hat{q}_{i}:=\log p_{\hat{z}_{i}|\mathbf{e},\mathbf{y}}$. Hence we have:
$\log p_{\mathbf{z}|\mathbf{e},\mathbf{y}}(\mathbf{z}|\mathbf{e},\mathbf{y})=\sum_{i=1}^{n}q_{i}(z_{i},\mathbf{e},\mathbf{y});\quad\log p_{\hat{\mathbf{z},\mathbf{y}}|\mathbf{e}}=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}).$ (13)
By combining Equation (13) and Equation (11), we have:
$p_{\mathbf{z}|\mathbf{e},\mathbf{y}}=p_{h(\hat{\mathbf{z}}|\mathbf{e},\mathbf{y})}\Longleftrightarrow p_{\hat{\mathbf{z}}|\mathbf{e},\mathbf{y}}=p_{\mathbf{z}|\mathbf{e},\mathbf{y}}|\mathbf{J}_{h^{-1}}|\Longleftrightarrow\sum_{i=1}^{n}q_{i}({z_{i},\mathbf{e}},\mathbf{y})+\log|\mathbf{J}_{h^{-1}}|=\sum_{i=1}^{n}\hat{q}_{i}(\hat{z}_{i},\mathbf{e},\mathbf{y}),$ (14)
where $\mathbf{J}_{h^{-1}}$ are the Jacobian matrix of $h^{-1}$.
Sequentially, we take the first-order derivative with $\hat{c}_{j}$ on Equation (14), where $j\in\{1,\cdots,n_{c}\}$, and have
$\sum_{i=1}^{n}\frac{\partial q_{i}(z_{i},\mathbf{u},\mathbf{y})}{\partial z_{i}}\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}+\frac{\partial\log|\mathbf{J}_{h^{-1}}|}{\partial\hat{c}_{j}}=\frac{\partial q_{j}(\hat{c}_{j},\mathbf{u},\mathbf{y})}{\partial\hat{c}_{j}}.$ (15)
Suppose $\mathbf{e}=e_{0},e_{1},\cdots,e_{n_{b}}$, we subtract the Equation (15) corresponding to $e_{k}$ with that corresponds to $e_{0}$, and we have:
$\sum_{i=1}^{n}\left(\frac{\partial q_{i}(z_{i},e_{k},\mathbf{y})}{\partial z_{i}}-\frac{\partial q_{i}(z_{i},e_{0},\mathbf{y})}{\partial z_{i}}\right)\cdot\frac{\partial z_{i}}{\partial\hat{c}_{j}}=\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}.$ (16)
Since the distribution of estimated $\hat{\mathbf{c}}_{j}$ does not change across different domains, $\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{k},\mathbf{y})}{\partial\hat{c}_{j}}-\frac{\partial\hat{q}_{j}(\hat{c}_{j},e_{0},\mathbf{y})}{\partial\hat{c}_{j}}=0$. Since $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial z_{i}}$ does not change across different domains, $\frac{\partial q_{i}(c_{i},e_{k},\mathbf{y})}{\partial c_{i}}=\frac{\partial q_{i}(c_{i},e_{0},\mathbf{y})}{\partial c_{i}}$ for $i\in\{1,\cdots,n_{c}\}$. So we have
$\sum_{i=1}^{n_{b}}\left(\frac{\partial q_{i}(b_{i},e_{k},\mathbf{y})}{\partial b_{i}}-\frac{\partial q_{i}(b_{i},e_{0},\mathbf{y})}{\partial b_{i}}\right)\cdot\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0.$ (17)
Based on the linear independence assumption (A3), the linear system is a full-rank system. Therefore, the only solution is $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$.
Since $h(\cdot)$ is smooth over $\mathbf{z}$, its Jacobian can be formalized as follows
$\begin{gathered}\mathbf{J}_{h}=\begin{bmatrix}\begin{array}[]{c|c}\textbf{A}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{b}}}&\textbf{B}:=\frac{\partial\mathbf{b}}{\partial\hat{\mathbf{c}}}\\ \hline\cr\textbf{C}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{b}}}&\textbf{D}:=\frac{\partial\mathbf{c}}{\partial\hat{\mathbf{c}}}.\end{array}\end{bmatrix}\end{gathered}$ (18)
Note that $\frac{\partial b_{i}}{\partial\hat{c}_{j}}=0$ for $i\in\{1,\cdots,n_{b}\}$ and $j\in\{1,\cdots,n_{c}\}$ means that $\textbf{B}=0$. Since $h(\cdot)$ is invertible, $\mathbf{J}_{h}$ is a full-rank matrix. Therefore, for each $b_{i},i\in\{1,\cdots,n_{b}\}$, there exists a $h_{i}$ such that $b_{i}=h_{i}(\hat{\mathbf{b}})$.
Besides, based on A4, one can show that all entries in the submatrix C zero according to part of the proof of Theorem 4.2 in [7](Steps 1, 2, and 3). Therefore, $\mathbf{b}$ and $\mathbf{c}$ are block-wise identifiable. ∎
B.2 Better predictor with adding bias
Proof.
Under Assumptions A1–A5, we compare the Bayes-optimal risks for predictors based on $c$ versus $(c,b)$.
Log-loss.
The Bayes-optimal log-loss risks satisfy
where
By A5 there is a positive-measure set of $(c,b)$ on which $p(y\mid c,b)\neq p(y\mid c)$. Hence
so $H(Y\mid C,B)<H(Y\mid C)$ and therefore
Squared-error loss.
The Bayes-optimal squared-error risks satisfy
By the law of total variance,
Since A5 ensures that $B$ influences the conditional mean of $Y$ given $C$, the second term is strictly positive on a set of positive probability. Thus
and consequently
This completes the proof that incorporating $b$ strictly reduces both log-loss and squared-error Bayes risks. ∎
@@ Appendix C Multi-class Method
Suppose we have $K\geq 2$ classes. We "one-hot encode" these classes, so that $y$ takes values in the set $\mathcal{Y}=\{(1,0,...,0),(0,1,0,...,0),...,(0,...,0,1)\}\subseteq\{0,1\}^{K}.$ In the training stage, let $\hat{y}=f_{\mathbf{c}}(\mathbf{c})$ $\epsilon\in[0,1]^{\mathcal{Y}\times\mathcal{Y}}$ with
@@
$\hat{y}\textbf{ only depends on }y$
@@
When $\epsilon$ is non-singular, this has the unique solution $E[y\mid\mathbf{b}]=\epsilon^{-1}E[\hat{y}\mid\mathbf{b}]$, giving a multiclass equivalent. However, it is numerically more stable to estimate $E[y\mid\mathbf{b}]$ by the least-squares solution
@@
Appendix D Proof of section 2
D.1 Decomposition and invariant
$=\frac{p(\mathbf{c},\mathbf{b},y=1)}{p(\mathbf{c},\mathbf{b},y=0)}$
$\forall e_{i},e_{j}\in e_{all}$ (23)
[…]
${p(\mathbf{c}\mid y=1,e=e_{i})}=p(\mathbf{c}\mid y=1,e=e_{j})$
   Use Bayes rule
D.2 Proof of theorem 2.5
In this section, we would like to prove $h_{0}+h_{1}=1$ if and only if $\mathbf{c}\perp\!\!\!\perp y$.
Forward Implication: Suppose $h_{0}+h_{1}=1$.
$h_{0}+h_{1}=1\quad\text{if and only if}\quad\mathbf{c}\perp y.$
Note that the entire result also holds, with almost identical proof, in the multi-environment setting of Sections 3 and 5, conditioned on a particular environment $E$.
Proof. We first prove the forward implication.
Suppose $h_{0}+h_{1}=1$.
If $\Pr[y=1]\in\{0,1\}$, then $\mathbf{c}$ and $y$ are trivially independent, because the value of $y$ is constant. In this case, the result holds trivially. So, we assume $\Pr[y=1]\in(0,1)$.
We now compute the expected value of the pseudo-label $\hat{y}$ using the Law of Total Expectation:
$E[\hat{y}]=\Pr[\hat{y}=1\mid y=1]\cdot\Pr[y=1]+\Pr[\hat{y}=0\mid y=0]\cdot\Pr[y=0].$
That is,
$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p),$
where $p=\Pr[y=1]$.
Next, we simplify this expression:
$E[\hat{y}]=h_{1}\cdot p+(1-h_{0})\cdot(1-p)=(h_{0}+h_{1}-1)\cdot p+1-h_{0}.$
Using the assumption $h_{0}+h_{1}=1$, we get:
$E[\hat{y}]=1-h_{0}.$
Thus, $E[\hat{y}]=1-h_{0}$.
Now, since the expected value of $\hat{y}$ is independent of the actual value of $y$, this suggests that $\hat{y}$ provides no information about $y$, implying that $\hat{y}$ and $y$ are independent.
Since $\hat{y}$ is a deterministic function of $\mathbf{c}$, this implies that $\mathbf{c}$ and $y$ must also be independent. Therefore, we have shown that if $h_{0}+h_{1}=1$, then $\mathbf{c}\perp y$.
To prove the reverse implication:
Suppose $\mathbf{c}\perp y$.
If $\mathbf{c}$ and $y$ are independent, then the pseudo-label $\hat{y}$ is also independent of $y$. Hence, we have:
$\Pr[\hat{y}=1\mid y=1]=\Pr[\hat{y}=1\mid y=0].$
This implies that:
$h_{1}=1-h_{0}.$
Therefore, we have:
$h_{0}+h_{1}=h_{0}+(1-h_{0})=1.$
Thus, we have shown that if $\mathbf{c}\perp y$, then $h_{0}+h_{1}=1$.
Putting the forward and reverse implications together, we conclude that:
$h_{0}+h_{1}=1\quad\Longleftrightarrow\quad\mathbf{c}\perp y.$
This completes the proof.
D.3 The derivation of the KL optimization
We want to solve the following problem:
Let
By the definition of KL divergence, for discrete $y$:
Hence, for each $e$,
Therefore,
To find the minimizing $q$, we set
[…]
First, note
Only the term $-\log q(y\mid{\mathbf{b}})$ depends on $q$. For a specific $y$,
Hence, for each $e$:
Summing over $e$,
Also,
Hence,
Setting this to zero:
$-\frac{1}{q(y\mid{\mathbf{b}})}\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e)+\lambda=0,$
which gives
Thus
We require $\sum_{y}q(y\mid{\mathbf{b}})=1$. Then:
Since $\sum_{y}p(y\mid{\mathbf{b}},e)=1$ for each $e$, we get
Hence
$11=\frac{1}{\lambda}\cdot 1\quad\Longrightarrow\quad\lambda=$ 1 .
Therefore,
$q(y\mid{\mathbf{b}})=\sum_{e}p(e\mid{\mathbf{b}})\,p(y\mid{\mathbf{b}},e).$
Appendix E Algorithms
E.1 Binary Bias-Aware Generalization
@@
2:  : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding, $Pr$ means logits prior, $\operatorname{logit}(p(y))$.
@@
4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
@@
7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
@@
16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
@@
E.2 Multi-Class Bias-Aware Generalization
@@
4:  Obtain $[\mathbf{c},\mathbf{b}]$ via feeding ${\mathbf{x}}$ into encoder.
@@
7:  Calculate the results of the invariant predictor $f_{c}{{\mathbf{c}}}$.
@@
16:  Fine-tune bias predictor $f_{b}$ with $(\mathbf{b}^{\mathcal{T}},\hat{y})$
@@
Appendix F Experiment details and ablation Study
F.1 Experiment details
@@
We evaluate our BAG model on the PACS dataset, which comprises four distinct domains (Photo, Art painting, Cartoon, Sketch) and seven object categories. Following the standard leave-one-domain-out protocol, we train on three domains and test on the held-out domain. Our backbone is a ResNet-18 pretrained on ImageNet, used to extract visual features. On top of this, we attach a variational autoencoder that separates causal factors from spurious factors, and we introduce learnable domain embeddings to encourage independence across environments. All key hyperparameters—such as the dimensions of the latent representations, the relative weighting between reconstruction and regularization losses, and the learning rates—are chosen via Bayesian optimization, with test-domain accuracy serving as the selection criterion. We train with the Adam optimizer for 70 epochs and select the final model based on its performance under test-time augmentation on the held-out domain. We used the code based on [11].
@@
We evaluate our BAG model on the Office-Home dataset [13], which comprises four domains (Art, Clipart, Product, Real-World) and 65 object categories. Following the leave-one-domain-out protocol, we train on three domains and test on the held-out one. Hyperparameters for each held-out domain are optimized via Bayesian optimization; the selected values for learning rate, regularization weight, number of epochs for the $z_{s}$ classifier, and VAE weight are given in Table F.1. The full code are provided in the supplementary material.
Table F.1: Optimal hyperparameters on Office-Home
Domain Learning Rate $\lambda_{1}$ epochs (test time) $\lambda_{0}$
@@
F.2 Ablation Study
Table F.2 presents the performance of various BAG variants on the PACS dataset and in simulation experiments. All variants outperform ERM, demonstrating the effectiveness of our two bias-utilization methods. Notably, BAG-RE and BAG-TTA each exhibit their own strengths and weaknesses, which may be due to data-specific effects. BAG-VAE performs somewhat worse than BAG, underscoring the importance of accurately identifying and disentangling $\mathbf{b}$ and $\mathbf{c}$.
Table F.2: Accuracy of different variants including BAG, BAG-VAE, BAG-RE, and BAG-TTA on four PACS domains and Synthetic data.
@@ F.2 Ablation Study
Appendix G Statement

### v2

@@ Appendix C Multi-class Method
Suppose we have $K\geq 2$ classes. We ”one-hot encode” these classes, so that $y$ takes values in the set $\mathcal{Y}=\{(1,0,...,0),(0,1,0,...,0),...,(0,...,0,1)\}\subseteq\{0,1\}^{K}.$ In the training stage, let $\hat{y}=f_{\mathbf{c}}(\mathbf{c})$ $\epsilon\in[0,1]^{\mathcal{Y}\times\mathcal{Y}}$ with
@@
y^ only depends on $\hat{y}\textbf${ only depends on }y
@@
When $\epsilon$ is non-singular, this has the unique solution $E[y\mid\mathbf{b}]=\epsilon^{-1}E[\hat{y}\mid\mathbf{b}]$, giving a multiclass equivalent. However, it is numerically more stable to estimate $E[y\mid\mathbf{b}]$ by the least-squares solution.
@@
Appendix D Algorithms
D.1 Binary Bias-Aware Generalization
@@
2: : In the following content $\mathcal{L}$ means loss, $f$ means classier, $e_{k}$ is the domain embedding, PrPr means logits prior, $\operatorname{logit}(p(y))$.
@@
4: Obtain $\mathbf{c},\mathbf{b}]$ via feeding $\mathbf{x}$} into encoder.
@@
7: Calculate the results of the invariant predictor $f_{c}{({\mathbf{c}})}$ .
@@
[…]
@@
D.2 Multi-Class Bias-Aware Generalization
@@
4: Obtain $\mathbf{c},\mathbf{b}]$ via feeding $\mathbf{x}$} into encoder.
@@
7: Calculate the results of the invariant predictor $f_{c}{({\mathbf{c}})}$.
@@
[…]
@@
Appendix E Experiment details and ablation Study
E.1 Experiment details
@@
We evaluate our BAG model on the PACS dataset, which comprises four distinct domains (Photo, Art painting, Cartoon, Sketch) and seven object categories. Following the standard leave-one-domain-out protocol, we train on three domains and test on the held-out domain. Our backbone is a ResNet-18 pretrained on ImageNet, used to extract visual features. On top of this, we attach a variational autoencoder that separates causal factors from spurious factors, and we introduce learnable domain embeddings to encourage independence across environments. All key hyperparameters—such as the dimensions of the latent representations, the relative weighting between reconstruction and regularization losses, and the learning rates—are chosen via Bayesian optimization, with test-domain accuracy serving as the selection criterion. We train with the Adam optimizer for 70 epochs and select the final model based on its performance under test-time augmentation on the held-out domain. We used the code based on (Jiang and Veitch, 2022).
@@
We evaluate our BAG model on the Office-Home dataset (Venkateswara et al., 2017), which comprises four domains (Art, Clipart, Product, Real-World) and 65 object categories. Following the leave-one-domain-out protocol, we train on three domains and test on the held-out one. Hyperparameters for each held-out domain are optimized via Bayesian optimization; the selected values for learning rate, regularization weight, number of epochs for the $z_{s}$ classifier, and VAE weight are given in Table 3. The full code are provided in the supplementary material.
Table 3: Optimal hyperparameters on Office-Home
Domain Learning Rate $\lambda_{1}$ # epochs (test time) $\lambda_{0}$
@@
Computing resources
We evaluate the BAG method on DomainNet under the leave-one-domain-out protocol, using clipart as the target domain and infograph, painting, quickdraw, real, and sketch as source domains, trained for 30 epochs with a fixed random seed and batch size 64. The experiment is conducted on a single AMD Instinct MI210 GPU (ROCm). Peak GPU memory usage reported by PyTorch is 26.43 GB, and the total wall-clock time is 21,547.08 s (approximately 5 h 59 m 7 s). For SFB under the same protocol and settings (clipart as the target domain, 30 epochs, batch size 64, fixed seed, single AMD Instinct MI210 GPU), the training time is 21,583.46 s (approximately 5 h 59 m 43 s) and the total time is 21,584.40 s (approximately 5 h 59 m 44 s), with a peak PyTorch-allocated memory of 26.35 GB. Compared with the most relevant baseline, BAG achieves lower runtime under similar memory usage; the peak memory is dominated by the ResNet-50 backbone and its training-time activations and optimizer states, which are shared across both methods, so the overall memory footprints are nearly identical.
E.2 Ablation Study
Table 4 presents a systematic ablation of BAG on PACS and in our synthetic simulations. The full method comprises two optimization stages and a VAE module. In the ablations, we remove exactly one component at a time, as indicated by a hyphen in the variant name: BAG-VAE drops the VAE in favor of a linear projection; BAG–TTA removes the Test Time Adaptation that uses pseudo labels to guide the bias branch; and BAG–RE removes the Reweight Experents mechanism. For transparency on the predictive roles of the two branches, Table 5 further reports the behavior of the bias predictor and the content predictor when used in isolation.
Empirically, all BAG variants outperform ERM across datasets, confirming the effectiveness of our bias–utilization design. BAG–RE and BAG–TTA show complementary strengths and weaknesses—consistent with data–specific factors such as domain composition and the severity of distribution shift—whereas BAG–VAE lags behind the full BAG, underscoring the value of accurately identifying and disentangling $\mathbf{b}$ (bias–aligned factors) and $\mathbf{c}$ (content–aligned factors). Using only the bias predictor or only the content predictor yields average accuracies of roughly $78\%$ and $80\%$, respectively; each alone lacks sufficient information and, without a learnable class prior, is vulnerable to label shift. Once the two predictors are combined under a learnable label prior, performance rises to state–of–the–art levels among the baselines considered.
In summary, the ablations show that each component contributes meaningfully: (i) explicit disentanglement of bias and content, (ii) expert reweighting and test–time adaptation to stabilize routing, and (iii) a compact generative prior via the VAE. Together, these elements enable BAG to deliver robust and consistently strong domain–generalization performance, while simplified variants remain competitive and illuminate the role of each design choice.
Table 4: Accuracy of different variants including BAG, BAG-VAE, BAG-RE, and BAG-TTA on four PACS domains and Synthetic data.
@@ E.2 Ablation Study
Table 5: Different parts performance results on PACS. Values are accuracies (%).
Method P A C S Avg
Content predictor Only 95.57 81.10 75.00 67.83 79.88
Bias predictor Only 94.01 73.58 78.67 66.23 78.12
BAG 96.29 83.69 80.29 76.05 84.08
E.3 Sensitivity of experts and embeddings
Table 6: Sensitivity with fixed experts E=3E{=}3 and varying embedding dimension $d$. Column-wise best in bold. Values are accuracies (%).
DsDs $P$ AA $C$ $S$ average
5 95.81 84.03 78.71 70.73 82.32
10 96.29 83.69 80.29 76.05 84.08
15 96.53 84.08 69.24 71.54 80.35
20 96.23 83.30 75.85 76.00 82.85
Table 7: Sensitivity with fixed embedding dimension d=10d{=}10 and varying number of experts $E$. Column-wise best in bold. Values are accuracies (%).
$E$ $P$ AA $C$ $S$ average
1 95.51 83.11 75.00 74.65 82.07
3 96.29 83.69 80.29 76.05 84.08
4 95.27 83.59 77.52 68.92 81.33
5 95.51 84.28 75.13 73.15 82.02
6 96.23 85.06 73.98 73.91 82.29
7 95.45 84.08 74.62 73.63 81.94
10 96.35 83.98 78.92 73.63 83.22
Is the environmental predictor necessary, or could one directly predict the label with similar performance? Yes, the environmental predictor is necessary. Directly predicting $y$ from bias–aligned features $b$ effectively fits $p(y\mid b)$, which is brittle under label shift: when $p(y)$ changes across domains, the conditional $p(y\mid b)$ drifts, and target performance degrades. Our approach explicitly estimates a latent environment and routes each example to domain–conditioned experts, separating environment–specific variation from task–relevant signals and improving robustness. As shown in Table 4, removing the environmental predictor (method BAG–RE) reduces the mean accuracy to $81.8\%$, below that of the full model. In our hyperparameter sweeps (Tables 6 and 7), most configurations that include the predictor exceed $81.8\%$ on average, and the best setting with E=3,d=10E{=}3,d{=}10 reaches $84.08\%$ with strong performance on the challenging domains $C{=}80.29\%$ and $S{=}76.05\%$. In summary, conditioning the classifier on an explicitly predicted environment addresses distribution shift and yields measurable, domain–wise robust gains over direct prediction.
How sensitive is performance to the number of experts and the embedding dimensionality? Sensitivity is non–monotonic in both factors. With experts fixed at E=3E{=}3 (Table 6), a medium environment embedding performs best: d=10d{=}10 attains the highest average $84.08\%$, whereas a smaller embedding underfits ($d{=}5:\ 82.32\%$). With embedding fixed at d=10d{=}10 (Table 7), increasing experts from E=1E{=}1 to E=3E{=}3 improves the average from $82.07\%$ to $\textbf{84.08}\%$; further increases provide no consistent gains (e.g., $E{=}10:\ 83.22\%$). A few well–specialized experts ($E{\approx}3$) together with a compact environment embedding ($d{\approx}10$) strike a reliable balance between expressivity and out–of–domain generalization. To summarize: BAG is not sensitive to these two parameters, and the performance is within a controllable range; even suboptimal parameter choices do not push it below baselines such as ERM, IRM and ACTIR.
E.4 Disentanglement of Latent Variables
We qualitatively examine the learned disentanglement on PACS by visualizing attributions for the content branch $\mathbf{c}$} and the bias branch $\mathbf{b}$} with Grad-CAM. Concretely, we reuse the trained checkpoint and the identical backbone as training. Given an input image $x$, we route it through the original BAG model but expose two thin wrappers that return the logits of the $\mathbf{c}$} head ($f_{\beta}$) and the $\mathbf{b}$} head ($f_{\eta}$), respectively. For Grad-CAM, we target the last residual block of layer4 in the backbone. We then compute two grayscale attribution maps: one for $\mathbf{c}$} by backpropagating through $f_{\beta}$, and one for $\mathbf{b}$} by backpropagating through $f_{\eta}$, each with ClassifierOutputTarget set to the ground-truth class.
To make the overlays visually comparable across images, we (i) bilinearly resize each Grad-CAM map to the input resolution; (ii) min–max normalize it to $[0,1]$; and (iii) render red intensity for $\mathbf{c}$} and blue intensity for $\mathbf{b}$} before alpha-blending with the RGB image (default $\alpha=0.5$). For each sample we compose a four-tile horizontal panel in the fixed order: original image, $\mathbf{b}$} (blue), $\mathbf{c}$} (red), and the composite overlay; the titles b/c are boldfaced to reduce ambiguity.
Across the shown examples, $\mathbf{c}$} consistently concentrates on class-relevant object regions, whereas $\mathbf{b}$} emphasizes background textures, styles, or peripheral cues. These observations align with the intended decomposition and support the use of $\mathbf{c}$} as content and $\mathbf{b}$} as environment/style bias in subsequent modules.
[…]
Appendix F Statement